# StereoFoley: Object-Aware Stereo Audio Generation from Video

### v3

@@ 4 Conclusion
We presented an end-to-end framework for video-to-audio generation. We showed that full-band stereo Foley generation can achieve state-of-the-art performance using SyncFormer with a simple latent-space matching strategy. More importantly, we demonstrated that the main challenge for object-aware stereo audio generation is not architectural but data-related. By synthesizing object-aware stereo datasets, the same model can be trained to produce spatially accurate, object-position–aware audio. While our pipeline uses amplitude panning, it could be adapted to alternative mixing conventions, extended to Ambisonics using object trajectories, or scaled to multichannel video-to-audio generation. Future work will explore these directions, alongside improving dataset quality and increasing diversity with both real and synthetic examples.

### v4

@@ 4 Conclusion
We presented an end-to-end framework for video-to-audio generation. We showed that our model achieves performance on par with state-of-the-art V2A models for full-band stereo audio generation, using a Synchformer video encoder with a simple latent-space matching strategy. More importantly, we demonstrated that the main challenge for object-aware stereo audio generation is not architectural but data-related. By synthesizing object-aware stereo datasets, the same model can be trained to produce spatially accurate, object-position–aware audio. While our pipeline uses amplitude panning, it could be adapted to alternative mixing conventions, extended to Ambisonics using object trajectories, or scaled to multichannel video-to-audio generation. Future work will explore these directions, alongside improving dataset quality and increasing diversity with both real and synthetic examples.